Given 2 quarters and 3 dimes how many different amounts of money can be created using 1 or more coins
Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

D: 10 cents
DD: 20
DDD: 30
Q: 25
QD: 35
QDD: 45
QDDD: 55
QQ: 50
QQD: 60
QQDD: 70
QQDDD: 80

Answer by ikleyn(53765)

.

how many = 3*4 - 1 = 12 - 1 = 11.         ANSWER

        Explanation



3  opportunity for quarters  (0, 1, 2)
and  4  opportunity for dimes  (0, 1, 2, 3).

So the product   3*4 = 12   is the number of all different possible combinations  (pairs  (q,d) ).


Minus the degenerated case  (q,d) = (0,0)  gives finally the answer   12-1 = 11.

--------------

Solved.